OpenStudy (anonymous):

solve this system of linear equations separate the x- and y-values with a comma 4x+7y=-3 8x+2y=-42

OpenStudy (anonymous):

d

OpenStudy (anonymous):

just solve for x or y in one of them and plug it into the other equation

OpenStudy (anonymous):

4x+2y=-42 2x+y=-21 y=-21-2x 4x+7y=-3 4x+7(-21-2x)=-3 and solve from there

OpenStudy (anonymous):

i still dont get it